i need help finding the slope of the line passing through the pointss (2,1),(6,9)?
Answer by haileytucki(390)

(2,1),(6,9)
Slope is equal to the change in y over the change in x, or 'rise over run'.
m=(change in y)/(change in x)
The change in x is equal to the difference in x-coordinates (also called run), and the change in y is equal to the difference in y-coordinates (also called rise).
m=(y2-y1)/(x2-x1)
Substitute in the values of x and y into the equation to find the slope.
m=(9-(1))/(6-(2))
Multiply -1 by the 2 inside the parentheses.
m=(9-(1))/(6-2)
Subtract 2 from 6 to get 4.
m=(9-(1))/(4)
Multiply -1 by the 1 inside the parentheses.
m=(9-1)/(4)
Subtract 1 from 9 to get 8.
m=(8)/(4)
Reduce the expression (8)/(4) by removing a factor of 4 from the numerator and denominator.
m=2